# Null-Space Filtering for Data-Free Continual Model Merging: Preserving Transparency, Promoting Fidelity

### v1

@@ Abstract
Data-free continual model merging (DFCMM) aims to fuse independently fine-tuned models into a single backbone that evolves with incoming tasks without accessing task data. This paper formulate two fundamental desiderata for DFCMM: transparency, avoiding interference with earlier tasks, and fidelity, adapting faithfully to each new task. This poses a challenge that existing approaches fail to address: how to bridge data-level desiderata with parameter-space optimization to ensure transparency and fidelity in the absence of task data. To this end, we propose NUFILT (NUll-space FILTering), a data-free framework that directly links these desiderata to optimization. Our key observation is that task vectors approximately align with representation subspaces, providing structural surrogates for enforcing transparency and fidelity. Accordingly, we design a null-space projector that preserves prior responses by filtering out overlapping components of new task vectors, thereby ensuring transparency, and a lightweight LoRA adapter that injects complementary task-specific signals, enabling fidelity in adapting to new tasks. The adapter is trained with a projection-based surrogate loss to retain consistency with previous knowledge while introducing novel directions. This joint filtering–adaptation process allows the backbone to absorb new knowledge while retaining existing behaviors, and the updates are finally fused back in a layer-wise linear fashion without extra parameters or inference cost. Theoretically, we establish approximate subspace alignment guarantees that justify null-space filtering. Empirically, NUFILT achieves state-of-the-art performance with minimal forgetting on both vision and NLP benchmarks, improving average accuracy by 4–7% over OPCM and WUDI-Merging, while narrowing the gap to fine-tuning and reducing computation overhead.
@@ 1 Introduction
Modern machine learning systems are often deployed in dynamic environments where tasks arrive one after another. Training a new model from scratch for each task is not only computationally expensive but also requires retaining all historical training data, which is often impractical in real applications. At the same time, maintaining a separate checkpoint for every task quickly becomes infeasible due to memory and deployment constraints (McMahan et al., 2017; Fang et al., 2024; Qiu et al., 2024). A more appealing alternative is to reuse existing models—either drawn from a model library or trained for new tasks—and consolidate their knowledge into a single backbone that evolves over time (Zhou et al., 2024b; Yu et al., 2024b; Huang et al., 2024a; Tang et al., 2025). However, this consolidation is expected to be performed while retaining only the merged model for storage efficiency, and without accessing any data in order to safeguard privacy, which makes the problem particularly challenging. These constraints highlight the urgent need for approaches that integrates knowledge across sequential tasks directly in parameter space without extra storage or data costs.
@@
Recent studies investigate data-free continual model merging (DFCMM) (Liu & Soatto, 2023; Porrello et al., 2025; Tang et al., 2025), where tasks arrive sequentially and, at each step, only the newly fine-tuned task model and the previously merged backbone are available (see Fig. 1). We formulate two fundamental desiderata for DFCMM: transparency, avoiding interference with earlier tasks, and fidelity, adapting faithfully to each new task. These desiderata, not explicitly considered in prior work, highlight a new open challenge for the field. Since they are inherently defined in terms of data, existing approaches lack a principled translation of data-level desiderata into parameter-space objectives. As a result, simple averaging or arithmetic updates often cause interference that undermines transparency (Izmailov et al., 2018; Ilharco et al., 2023); projection methods fail to preserve fidelity when task vectors are correlated (Tang et al., 2025; Yadav et al., 2023); and adaptive strategies typically rely on auxiliary data, which violates the data-free constraint (Yang et al., ; Tang et al., 2024b; Qiu et al., 2025). Achieving both transparency and fidelity under the strict requirements of DFCMM therefore remains an open and unresolved challenge.
@@
Figure 1: Illustration of data-free continual model merging (DFCMM). At each step, only the current task model and the previously merged model are accessible, and the merging process is performed without access to any data. The merged model is expected to preserve prior knowledge (transparency) while adapting efficiently to new tasks (fidelity).
@@
In this paper, we introduce NUFILT (NUll-space FILTering), a novel framework for DFCMM that directly bridges data-level desiderata with parameter-space optimization. The key observation is that task vectors exhibit approximate alignment with representation subspaces, enabling structural surrogates for enforcing transparency and fidelity without data. To achieve transparency, a null-space projector filters out components of the new task vector overlapping with earlier subspaces, suppressing interfering activations. To achieve fidelity, a lightweight LoRA adapter performs projection-aware adaptation, optimized with a data-free surrogate loss: updates are constrained to remain consistent with previous tasks while introducing complementary signals along directions unique to the new task. Finally, the projector, task vector, and adapter are fused back into the backbone in a layer-wise linear fashion, so updates are absorbed without extra parameters or inference cost. Theoretically, we provide guarantees of approximate subspace alignment, establishing a rigorous foundation for null-space filtering. Empirically, we demonstrate that NUFILT achieves state-of-the-art accuracy with minimal forgetting across both vision and NLP tasks. Compared to recent methods such as OPCM (Tang et al., 2025) and WUDI-Merging (Cheng et al., ), NUFILT improves average accuracy by 4–7%, while substantially narrowing the gap to individual fine-tuning and reducing computation overhead.
@@
  • We formulate transparency and fidelity as two fundamental desiderata for data-free continual model merging, framing a new open challenge absent from prior work.
@@
  • We propose NUFILT, a data-free continual model merging framework that combines null-space filtering with projection-aware adaptation to enforce both transparency and fidelity.
@@ 2 Related Work
Model Merging.  Model merging provides an efficient alternative to multi-task or continual training by combining multiple fine-tuned models directly in parameter space, without revisiting their training data. Early efforts adopted simple weight averaging (Utans, 1996; Shoemake, 1985), which can connect models through linear mode connectivity (Entezari et al., 2021; Ainsworth et al., 2022) but often suffers from severe performance degradation due to parameter interference. Task Arithmetic (TA) (Ilharco et al., 2023) formalizes merging through task vectors, making merging operations more explicit. However, since standard fine-tuning rarely guarantees disentangled updates (Ortiz-Jimenez et al., 2023), TA can amplify interference, prompting structured finetuning strategies (Ortiz-Jimenez et al., 2023; Liu & Soatto, 2023; Porrello et al., 2025). To mitigate conflicts, several methods introduce additional structure. Sparsity-based approaches such as TIES-Merging (Yadav et al., 2023) prune redundant or conflicting parameters, while reweighting schemes (Yu et al., 2024b) balance task contributions. Projection-based methods enforce orthogonality between task vectors to separate task directions (Tang et al., 2025; Cheng et al., ), but struggled when tasks are inherently correlated. Adaptive strategies further adjust merging process with auxiliary calibration data (Yang et al., ; Tang et al., 2025; Qiu et al., 2025), but this departs from the strictly data-free setting. Overall, these methods reduce interference to varying degrees but either rely on data or assume strong independence between tasks, leaving the core challenge of simultaneously achieving transparency and fidelity in the data-free continual setting largely unresolved.
@@
Continual Learning.  Continual learning tackles catastrophic forgetting (McCloskey & Cohen, 1989), where models lose performance on earlier tasks when adapting to new ones. Solutions include constraining parameter updates via importance weights (Kirkpatrick et al., 2016; Zenke et al., 2017; Aljundi et al., 2018), preserving knowledge through distillation (Hou et al., 2019; Douillard et al., 2020), or replaying exemplars and surrogates (Rebuffi et al., 2016; Liu et al., 2021). Other strategies expand capacity with dynamic architectures (Lee et al., 2017; Zhou et al., 2024a) or adopt parameter-efficient modules such as adapters and prompts (Yu et al., 2024a; Huang et al., 2024b). Beyond traditional data-driven approaches, a complementary direction explores continual learning through model merging. Early efforts merged fine-tuned checkpoints to alleviate forgetting (Mirzadeh et al., 2021; Wen et al., 2023; Marczak et al., 2024), but these typically relied on training data or sequential fine-tuning. More recently, continual model merging (Jin et al., 2023; Liu & Soatto, 2023; Porrello et al., 2025; Tang et al., 2025; Qiu et al., 2025) has emerged, aiming to fuse independently fine-tuned models directly in parameter space, avoiding both data access and checkpoint storage while improving scalability and privacy.
@@ 3.3 Desiderata for Data-Free Continual Merging
We identify two desiderata for the transformed task vectors $\tilde{\tau}_{i}=F(\tau_{i})$ when merging without data:
①Transparency to previous tasks. New task updates should not overwrite knowledge from earlier tasks. For any input $X$ from $\mathcal{D}_{i\leq t-1}$, the predictions after adding $\tilde{\tau}_{t}$ should remain consistent with those before:
$\mathcal{L}_{\text{trans}}=\mathbb{E}_{X\sim\mathcal{D}_{i\leq t-1}}\Big[\ell\big((\theta_{0}+\sum_{i=1}^{t}\tilde{\tau}_{i})X^{\top},\,(\theta_{0}+\sum_{i=1}^{t-1}\tilde{\tau}_{i})X^{\top}\big)\Big],$ (3)
where $\ell$ is a distance metric (e.g., squared $\ell_{2}$, KL divergence).
②Fidelity to the current task. The merged model should also reproduce the behavior of the task-specific model $\theta_{t}$ on its in-distribution inputs:
$\mathcal{L}_{\text{fid}}=\mathbb{E}_{X\sim\mathcal{D}_{t}}\Big[\ell\big((\theta_{0}+\sum_{i=1}^{t}\tilde{\tau}_{i})X^{\top},\,(\theta_{0}+\tau_{t})X^{\top}\big)\Big].$ (4)
In principle, optimizing $\mathcal{L}_{\text{trans}}$ and $\mathcal{L}_{\text{fid}}$ would yield task vectors that preserve prior knowledge while faithfully incorporating new tasks. Yet this is infeasible in the data-free setting, since $\mathcal{D}_{i\leq t-1}$ and $\mathcal{D}_{t}$ are unavailable. This leads to the central challenge:
How can continual merging achieve transparency and fidelity without data?
@@
Figure 2: Subspace affinity between data and task vectors in ViT-B/16 across eight datasets. Heatmaps show diagonal dominance, and layer-wise ECDFs confirm higher affinities for matched pairs.
@@ 3.4 Approximate Alignment Between Representation and Task Vector Subspaces
For each task and layer $l$, let $V_{d}^{(l)}\in\mathbb{R}^{d_{i}\times r_{d}}$ be the singular vectors of the representation covariance. For the task vector $\tau^{(l)}$, we compute its truncated SVD, $\tau^{(l)}\approx\hat{U}^{(l)}\hat{\Sigma}^{(l)}\hat{V}^{(l)\top}$, where $\hat{V}^{(l)}\in\mathbb{R}^{d_{i}\times r_{v}}$ is the top $r_{v}$ right singular vectors. We define the subspace affinity:
@@
Let $\tau^{(l)}\in\mathbb{R}^{d_{o}\times d_{i}}$ be the task vector at layer $l$, and $H\in\mathbb{R}^{N\times d_{i}}$ a representation matrix of rank $r_{d}$ with right singular vectors $V_{d}\in\mathbb{R}^{d_{i}\times r_{d}}$. Suppose $\tau^{(l)}$ admits the decomposition $\tau^{(l)}=T_{0}+E$ with $T_{0}=BH,\;\mathrm{rank}(T_{0})=r_{d}$, where each row $E_{k}$ of $E$ satisfies $\|E_{k}\|_{2}\leq\Psi_{k}$. If $\sigma_{r_{d}}(T_{0})>\sqrt{d_{o}}\,\max_{k}\Psi_{k},$ then for any $r_{v}\geq r_{d}$, the span of the top $r_{v}$ right singular vectors $\hat{V}$ of $\tau^{(l)}$ is approximately aligned with $\mathrm{span}(V_{d})$ in the sense that
$\mathcal{A}(V_{d}^{(l)},\hat{V}^{(l)})\;\leq\;\zeta^{2},$ (6)
where $\mathcal{A}(V_{d}^{(l)},\hat{V}^{(l)})=1-\tfrac{1}{r_{d}}\|\hat{V}^{\top}V_{d}\|_{F}^{2}$ and $\zeta^{2}=\left(\frac{\sqrt{d_{o}}\,\max_{k}\Psi_{k}}{\sigma_{r_{d}}(T_{0})-\sqrt{d_{o}}\,\max_{k}\Psi_{k}}\right)^{2}.$

### v2

# Null-Space Filtering for Data-Free Continual Model Merging: Preserving Stability, Promoting Plasticity

@@ Abstract
Data-free continual model merging (DFCMM) aims to fuse independently fine-tuned models into a single backbone that evolves with incoming tasks without accessing task data. This paper revisits two fundamental desiderata for DFCMM: stability, avoiding interference with earlier tasks, and plasticity, adapting faithfully to each new task. This poses a challenge that existing approaches fail to address: how to bridge data-level desiderata with parameter-space optimization to ensure stability and plasticity in the absence of task data. To this end, we propose NUFILT (NUll-space FILTering), a data-free framework that directly links these desiderata into parameter-space optimization. Our key observation is that task vectors approximately align with representation subspaces, providing structural surrogates for enforcing stability and plasticity. Accordingly, we design a null-space projector that preserves prior responses by filtering overlapping components of new task vectors, ensuring stability. We further introduce a lightweight LoRA adapter that injects complementary task-specific signals to enable plasticity. The adapter is trained with a projection-based surrogate loss that preserves consistency with prior knowledge while introducing novel directions. This joint filtering–adaptation process enables the backbone to absorb new knowledge while retaining existing behaviors, with updates fused back in a layer-wise linear fashion without extra parameters or inference cost. Theoretically, we establish approximate subspace alignment guarantees that justify null-space filtering. Empirically, NUFILT achieves state-of-the-art performance with minimal forgetting on both vision and NLP benchmarks, improving average accuracy by 4–7% over OPCM and WUDI-Merging, while narrowing the gap to fine-tuning and reducing computation overhead. The code is available at: https://github.com/zihuanqiu/NUFILT
@@ 1 Introduction
Modern machine learning systems are often deployed in dynamic environments where tasks arrive one after another. Training a new model from scratch for each task is computationally expensive and requires retaining all historical data, which is often impractical. Maintaining a separate checkpoint for every task also becomes infeasible due to memory and deployment constraints (McMahan et al., 2017; Fang et al., 2024; Qiu et al., 2024). A more appealing alternative is to reuse existing models—either drawn from a model library or trained for new tasks—and consolidate their knowledge into a single backbone that evolves over time (Zhou et al., 2024b; Yu et al., 2024b; Huang et al., 2024a; Tang et al., 2025). However, this consolidation is expected to be performed while retaining only the merged model for storage efficiency, and without accessing any data in order to safeguard privacy, which makes the problem particularly challenging. These constraints highlight the urgent need for approaches that integrate knowledge across sequential tasks directly in parameter space without extra storage or data costs.
@@
Recent studies investigate data-free continual model merging (DFCMM) (Liu and Soatto, 2023; Porrello et al., 2025; Tang et al., 2025), where tasks arrive sequentially and, at each step, only the newly fine-tuned task model and the previously merged backbone are available (see Fig. 1). We revisit two fundamental desiderata for DFCMM: stability, avoiding interference with earlier tasks, and plasticity, adapting faithfully to each new task. While these desiderata are inherently defined at the data level, existing approaches lack a principled translation into parameter-space objectives. As a result, simple averaging or arithmetic updates often cause interference that undermines stability (Izmailov et al., 2018; Ilharco et al., 2023); projection methods fail to preserve plasticity when task vectors are correlated (Tang et al., 2025; Yadav et al., 2023); and adaptive strategies typically rely on auxiliary data, which violates the data-free constraint (Yang et al., ; Tang et al., 2024b; Qiu et al., 2025). Achieving both stability and plasticity under the strict requirements of DFCMM remains an open challenge.
@@
Figure 1: Illustration of data-free continual model merging (DFCMM). At each step, only the current task model and the previously merged model are accessible, and the merging process is performed without access to any data. The merged model is expected to preserve prior knowledge (stability) while adapting efficiently to new tasks (plasticity).
@@
In this paper, we introduce NUFILT (NUll-space FILTering), a novel framework for DFCMM that directly bridges data-level desiderata with parameter-space optimization. The key observation is that task vectors exhibit approximate alignment with representation subspaces, enabling structural surrogates for enforcing stability and plasticity without data. To achieve stability, a null-space projector filters out components of the new task vector overlapping with earlier subspaces, suppressing interfering activations. To achieve plasticity, a lightweight LoRA adapter performs projection-aware adaptation, optimized with a data-free surrogate loss: updates are constrained to remain consistent with previous tasks while introducing complementary signals along directions unique to the new task. Finally, the projector, task vector, and adapter are fused back into the backbone in a layer-wise linear fashion, so updates are absorbed without extra parameters or inference cost. Theoretically, we provide guarantees of approximate subspace alignment, establishing a rigorous foundation for null-space filtering. Empirically, we demonstrate that NUFILT achieves state-of-the-art accuracy with minimal forgetting across both vision and NLP tasks. Compared to recent methods such as OPCM (Tang et al., 2025) and WUDI-Merging (Cheng et al., 2025), NUFILT improves average accuracy by 4–7%, while substantially narrowing the gap to individual fine-tuning and reducing computation overhead.
@@
  • We propose NUFILT, a data-free continual model merging framework that combines null-space filtering with projection-aware adaptation to enforce both stability and plasticity.
@@ 2 Related Work
Model Merging.  Model merging provides an efficient alternative to multi-task or continual training by combining multiple fine-tuned models directly in parameter space, without revisiting their training data. Early efforts adopted simple weight averaging (Utans, 1996; Shoemake, 1985), which can connect models through linear mode connectivity (Entezari et al., 2021; Ainsworth et al., 2022) but often suffers from severe performance degradation due to parameter interference. Task Arithmetic (TA) (Ilharco et al., 2023) formalizes merging through task vectors, making merging operations more explicit. However, since standard fine-tuning rarely guarantees disentangled updates (Ortiz-Jimenez et al., 2023), TA can amplify interference, prompting structured finetuning strategies (Ortiz-Jimenez et al., 2023; Liu and Soatto, 2023; Porrello et al., 2025). To mitigate conflicts, several methods introduce additional structure. Sparsity-based approaches such as TIES-Merging (Yadav et al., 2023) prune redundant or conflicting parameters, while reweighting schemes (Yu et al., 2024b) balance task contributions. Projection-based methods enforce orthogonality between task vectors to separate task directions (Tang et al., 2025; Cheng et al., 2025), but struggled when tasks are inherently correlated. Adaptive strategies further adjust merging process with auxiliary calibration data (Yang et al., ; Tang et al., 2025; Qiu et al., 2025), but this departs from the strictly data-free setting. Overall, these methods reduce interference to varying degrees but either rely on data or assume strong independence between tasks, leaving the core challenge of simultaneously achieving stability and plasticity in the data-free continual setting largely unresolved.
@@
Continual Learning.  Continual learning tackles catastrophic forgetting (McCloskey and Cohen, 1989), where models lose performance on earlier tasks when adapting to new ones. Solutions include constraining parameter updates via importance weights (Kirkpatrick et al., 2016; Zenke et al., 2017; Aljundi et al., 2018), preserving knowledge through distillation (Hou et al., 2019; Douillard et al., 2020), or replaying exemplars and surrogates (Rebuffi et al., 2016; Liu et al., 2021). Other strategies expand capacity with dynamic architectures (Lee et al., 2017; Qiu et al., 2023; Zhou et al., 2024a) or adopt parameter-efficient modules such as adapters and prompts (Yu et al., 2024a; Huang et al., 2024b). Beyond traditional data-driven approaches, a complementary direction explores continual learning through model merging. Early efforts merged fine-tuned checkpoints to alleviate forgetting (Mirzadeh et al., 2021; Wen et al., 2023; Marczak et al., 2024), but these typically relied on training data or sequential fine-tuning. More recently, continual model merging (Jin et al., 2023; Liu and Soatto, 2023; Porrello et al., 2025; Tang et al., 2025; Qiu et al., 2025) has emerged, aiming to fuse independently fine-tuned models directly in parameter space, avoiding both data access and checkpoint storage while improving scalability and privacy.
Orthogonality in Task Fusion.  Orthogonality-based methods aim to reduce interference between tasks by projecting updates into orthogonal spaces. These can be categorized into two types: orthogonality between data and parameters (Farajtabar et al., 2020; Wang et al., 2021; Liang and Li, 2024), and orthogonality directly between parameter updates(Wei et al., ; Cheng et al., 2025; Tang et al., 2025). Fang et al. (2025) project perturbations onto the null space of preserved knowledge, while Xiong et al. (2024) enforce orthogonality between task vectors, and Wang et al. (2023) learn new parameters within orthogonal low-rank subspaces to reduce interference. In contrast, our method projects each new task vector into the null space of previously merged task representations, but without requiring any task data. This differs from Fang et al. (2025); Wang et al. (2023), which rely on external data to estimate orthogonality constraints. Moreover, unlike Xiong et al. (2024), which requires simultaneous access to all task vectors, our approach operates sequentially using only the previously merged task vector and the current one, enabling efficient continual model merging.
@@ 3.3 Desiderata for Data-Free Continual Merging
In continual model merging, updates for a new task should be incorporated without disrupting knowledge accumulated from earlier ones. This reflects the classical stability–plasticity balance in continual learning. To describe how a transformed update $\tilde{\tau}^{(l)}_{t}=F(\tau^{(l)}_{t})$ influences the merged model, we use two layer-wise quantities as surrogates for these effects. Let $\theta^{(l)}_{0}\in\mathbb{R}^{d_{o}\times d_{i}}$ be the pretrained weights at layer $l$, $\tau^{(l)}_{i}$ the task vector for task $i$, and $x^{(l)}\in\mathbb{R}^{d_{i}}$ the corresponding layer activations.
①Stability surrogate. This term measures how much the new update alters the model’s responses on earlier tasks. For activations $x^{(l)}\sim\mathcal{D}_{i\leq t-1}$, we define:
$\mathcal{L}^{(l)}_{\text{stab}}(\tilde{\tau}^{(l)}_{t})=\mathbb{E}_{x^{(l)}\sim\mathcal{D}_{i\leq t-1}}\big[\ell((\theta_{t-1}^{\mathrm{merged},(l)}+\tilde{\tau}^{(l)}_{t})x^{(l)},\theta_{t-1}^{\mathrm{merged},(l)}x^{(l)})\big].$ (3)
where $\ell$ is a distance metric (e.g., squared $\ell_{2}$).
②Plasticity surrogate. This term evaluates how well the merged parameters follow the behavior of the task’s individual model. For activations $x^{(l)}\sim\mathcal{D}_{t}$, we define:
$\mathcal{L}^{(l)}_{\text{plas}}(\tilde{\tau}^{(l)}_{t})=\mathbb{E}_{x^{(l)}\sim\mathcal{D}_{t}}\big[\ell((\theta_{t-1}^{\mathrm{merged},(l)}+\tilde{\tau}^{(l)}_{t})x^{(l)},(\theta^{(l)}_{0}+\tau^{(l)}_{t})x^{(l)})\big].$ (4)
These two terms anchor the stability–plasticity trade-off within the parameter space used for merging. However, optimizing $\mathcal{L}^{(l)}_{\text{stab}}(\tilde{\tau}^{(l)}_{t})$ and $\mathcal{L}^{(l)}_{\text{plas}}(\tilde{\tau}^{(l)}_{t})$ is not possible in the data-free setting, as neither $\mathcal{D}_{i\leq t-1}$ nor $\mathcal{D}_{t}$ is accessible. This leads to the central challenge:
How can continual merging achieve stability and plasticity without data?
@@
Figure 2: Subspace affinity between data and task vectors in ViT-B/16 across eight datasets. Heatmaps show diagonal dominance, and layer-wise empirical cumulative distribution functions (ECDFs) confirm higher affinities for matched pairs.
@@ 3.4 Approximate Alignment Between Representation and Task Vector Subspaces
For each task and layer $l$, let $V_{d}^{(l)}\in\mathbb{R}^{d_{i}\times r_{d}}$ be the singular vectors of the representation covariance $\mathbb{E}[x^{(l)}(x^{(l)})^{\top}]$. For the task vector $\tau^{(l)}$, we compute its truncated SVD, $\tau^{(l)}\approx\hat{U}^{(l)}\hat{\Sigma}^{(l)}\hat{V}^{(l)\top}$, where $\hat{V}^{(l)}\in\mathbb{R}^{d_{i}\times r_{v}}$ is the top $r_{v}$ right singular vectors. We define the subspace affinity:
@@
Let $\tau^{(l)}\in\mathbb{R}^{d_{o}\times d_{i}}$ be the task vector at layer $l$, and $H\in\mathbb{R}^{N\times d_{i}}$ a representation matrix of rank $r_{d}$ with right singular vectors $V_{d}\in\mathbb{R}^{d_{i}\times r_{d}}$. Suppose $\tau^{(l)}$ admits the decomposition $\tau^{(l)}=T_{0}+E$ with $T_{0}=BH,\;\mathrm{rank}(T_{0})=r_{d}$, where each row $E_{k}$ of $E$ satisfies $\|E_{k}\|_{2}\leq\Psi_{k}$. Let $\sigma_{r_{d}}(T_{0})$ denote the $r_{d}$-th largest singular value of the matrix $T_{0}$. If $\sigma_{r_{d}}(T_{0})>\sqrt{d_{o}}\,\max_{k}\Psi_{k},$ then for any $r_{v}\geq r_{d}$, the span of the top $r_{v}$ right singular vectors $\hat{V}$ of $\tau^{(l)}$ is approximately aligned with $\mathrm{span}(V_{d})$ in the sense that
$1-\mathcal{A}(V_{d}^{(l)},\hat{V}^{(l)})\;\leq\;\zeta^{2},$ (6)
where $\mathcal{A}(V_{d}^{(l)},\hat{V}^{(l)})=\tfrac{1}{r_{d}}\|\hat{V}^{\top}V_{d}\|_{F}^{2}$ and $\zeta=\frac{\sqrt{d_{o}}\,\max_{k}\Psi_{k}}{\sigma_{r_{d}}(T_{0})-\sqrt{d_{o}}\,\max_{k}\Psi_{k}}.$